OpenStudy (huemath):

y=-5x +10 solve with out graphing? pls help

OpenStudy (t-dawg02):

you will probably get more help in the mathematics section http://openstudy.com/study#/groups/Mathematics and watch this awesome OS tutorial https://www.youtube.com/watch?v=lDiiH8PItto also when someone helps you click best response

ILovePuppiesLol (ilovepuppieslol):

Hello There! Welcome to Openstudy!! Since you are new here, read this tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/56f5205ce4b07a8c82287f75 This is the open study feedback section and you would likely to get more help in the Math section! http://openstudy.com/study#/groups/Mathematics

likeabossssssss (likeabossssssss):

welcome to os!!!!!!!!!! we r glad u r here

OpenStudy (fortytherapper):

I would suggest by first plotting points. Take some certain x values. For example let's do x = 1 (1, y) to find y, we plug in x y = -5(1) + 10 That means y = 5 (1,5) is a point you can plot. You can do this with any x value since it's a polynomial. I would keep the values low though.

TheSmartOne (thesmartone):

There must be another equation if you are trying to solve the system of equations.